OpenStudy (nuccioreggie):

A candy bar box is in the shape of a triangular prism. The volume of the box is 3,240 cubic centimeters. A triangular prism is shown with base of triangle labeled 24 cm, sides of the triangle labeled 15 cm, and length of the box equal to 30 cm. Part A: What is the height of the box? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make this candy box? Explain how you got your answer. (5 points)

OpenStudy (nuccioreggie):

Please help me and explain the answer

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (here_to_help15):

lol your doing this on android ?

OpenStudy (nuccioreggie):

Lol yea lololl

OpenStudy (nuccioreggie):

@Michele_Laino. Can you explain the answer

OpenStudy (alekos):

VOLUME = 1/2 base x height x length

OpenStudy (michele_laino):

the area of the base of your prism, is the area of a triangle whose base and height are 24cm and 15 cm respectively. SO the area of the base of your prism is: A=(24*15)/2=...cm^2

OpenStudy (michele_laino):

So*

OpenStudy (nuccioreggie):

Thats the answer for a and b ?

OpenStudy (michele_laino):

it is the first step for part A

OpenStudy (nuccioreggie):

Whats B?

OpenStudy (nuccioreggie):

Ok whats the next step for part A

OpenStudy (michele_laino):

what is the base area? please continue my computation

OpenStudy (nuccioreggie):

Ok so wait is is 180^2cm ?? Right

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now we can write: volume = base area x height of prism so, we get: \[3240 = 180 \times H\] what is H?

OpenStudy (nuccioreggie):

24cm

OpenStudy (michele_laino):

are you sure? 24*180=4320

OpenStudy (nuccioreggie):

15cm?

OpenStudy (michele_laino):

what is: 3240/180=...?

OpenStudy (nuccioreggie):

180 @Michele_Laino

OpenStudy (nuccioreggie):

@I ment its 18

OpenStudy (michele_laino):

ok! So the height of your prism is H=18 cm

OpenStudy (nuccioreggie):

Ok so i have to do 180x18?

OpenStudy (michele_laino):